# The failure of Hölder regularity of solutions for the Camassa–Holm type equation in Besov spaces

### 20 Jan 2024 v1

@@
Abstract: It is proved that if $u_{0}\in B^{s}_{p,r}$ with $s>1+\frac{1}{p},(p,r)\in[1,+\infty]\times[1,+\infty)$ or $s=1+\frac{1}{p},\ (p,r)\in[1,+\infty)\times\{1\}$, the solution of the Camassa–Holm equation belongs to $\mathcal{C}([0,T];B^{s}_{p,r})$. In the paper, we show that the continuity of the solution can not be improved to the Hölder continuity. Precisely speaking, the solution of the Camassa–Holm equation belongs to $\mathcal{C}([0,T];B^{s}_{p,r})$ but not to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with any $\alpha\in(0,1)$.
Keywords: Camassa–Holm equation, Hölder regularity, Besov spaces
MSC (2010): 35Q35; 35B30.
@@ 1 Introduction
In this paper, we are concerned with the Cauchy problem for the classical Camassa–Holm (CH) equation
@@
Here the scalar function $u=u(t,x)$ stands for the fluid velocity at time $t\geq 0$ in the $x$ direction.
@@
Setting $\Lambda^{-2}=(1-\partial^{2}_{xx})^{-1}$, then $\Lambda^{-2}f=G*f$ where $G(x)=\frac{1}{2}e^{-|x|}$ is the kernel of the operator $\Lambda^{-2}$. Thus, we can transform the CH equation (1.1) equivalently into the following transport type equation
@@
$\begin{cases}\partial_{t}u+u\partial_{x}u=\mathbf{P}(u),\;&(x,t)\in\mathbb{R}\times\mathbb{R}^{+},\\ u(x,t=0)=u_{0}(x),\;&x\in\mathbb{R},\end{cases}$ (1.2)
where
$\mathbf{P}(u)=P(D)\left(u^{2}+\frac{1}{2}(\partial_{x}u)^{2}\right)\quad\text{with}\quad P(D)=-\partial_{x}\Lambda^{-2}.$ (1.3)
@@
The CH equation (1.1) was firstly proposed in the context of hereditary symmetries studied in [16] and then was derived explicitly as a water wave equation by Camassa–Holm [2]. Many aspects of the mathematical beauty of the CH equation have been exposed over the last two decades. Particularly, (CH) is completely integrable [2, 5] with a bi-Hamiltonian structure [4, 16] and infinitely many conservation laws [2, 16]. Also, it admits exact peaked soliton solutions (peakons) of the form $u(x,t)=ce^{-|x-ct|}\;(c>0)$, which are orbitally stable [12]. Another remarkable feature of the CH equation is the wave breaking phenomena: the solution remains bounded while its slope becomes unbounded in finite time [3, 8, 9]. It is worth mentioning that the peaked solitons present the characteristic for the travelling water waves of greatest height and largest amplitude and arise as solutions to the free-boundary problem for incompressible Euler equations over a flat bed, see Refs. [6, 10, 11, 39] for the details.
Due to these interesting and remarkable features, the CH equation has attracted much attention as a class of integrable shallow water wave equations in recent twenty years. Its systematic mathematical study was initiated in a series of papers by Constantin and Escher, see [7, 8, 9, 10, 11]. After the CH equation was derived physically in the context of water waves, there are a large amount of literatures devoted to studying the well-posedness of the Cauchy problem (1.1) (see Molinet’s survey [37]). Li and Olver [36] proved that the Cauchy problem (1.1) is locally well-posed with the initial data $u_{0}\in H^{s}(\mathbb{R})$ with $s>3/2$ (see also [38]). Danchin [13, 14] proved the local existence and uniqueness of strong solutions to (1.1) with initial data in $B^{s}_{p,r}$ if $(p,r)\in[1,\infty]\times[1,\infty),s>\max\big{\{}1+1/p,3/2\big{\}}$ and $B^{3/2}_{2,1}$. Meanwhile, he [13] only obtained the continuity of the solution map of (1.1) with respect to the initial data in the space $\mathcal{C}([0,T];B^{s^{\prime}}_{p,r})$ with any $s^{\prime}<s$. Li-Yin [32] proved the continuity of the solution map of (1.1) with respect to the initial data in the space $\mathcal{C}([0,T];B^{s}_{p,r})$ with $r<\infty$. In particular, they [32] proved that the solution map of (1.1) is weak continuous with respect to initial data $u_{0}\in B^{s}_{p,\infty}$. For the endpoints, Danchin [14] obtained that the data-to-solution map is not continuous by using peakon solution, which implies the ill-posedness of (1.1) in $B^{3/2}_{2,\infty}$. Himonas-Misiołek [23] obtained the first result on the non-uniform dependence for (1.1) in $H^{s}(\mathbb{T})$ with $s\geq 2$ using explicitly constructed travelling wave solutions, which was sharpened to $s>\frac{3}{2}$ by Himonas-Kenig [21] on the real-line and Himonas-Kenig-Misiołek [22] on the circle. We should mention that, non-uniform continuity of the CH solution map in $H^{1}(\mathbb{R}\;\text{or}\;\mathbb{T})$ was established by Himonas-Misiołek-Ponce [24] by using traveling wave solutions. In our recent papers [31, 30], we proved the non-uniform dependence on initial data for (1.1) under both the framework of Besov spaces $B^{s}_{p,r}$ for $s>\max\big{\{}1+1/p,3/2\big{\}}$ with $(p,r)\in[1,\infty]\times[1,\infty)$ and $B^{3/2}_{2,1}$. Guo-Liu-Molinet-Yin [17] showed the ill-posedness of (1.1) in $B_{p,r}^{1+1/p}(\mathbb{R}\;\text{or}\;\mathbb{T})$ with $(p,r)\in[1,\infty]\times(1,\infty]$ (especially in $H^{3/2}$) by proving the norm inflation. Very recently, Guo-Ye-Yin [18] obtained the ill-posedness for the CH equation in $B^{1}_{\infty,1}(\mathbb{R})$ by proving the norm inflation. Ye-Yin-Guo [40] obtained the local well-posedness for the Camassa–Holm type equation in Besov spaces $s>1+\frac{1}{p},(p,r)\in[1,+\infty]\times[1,+\infty)$ or $s=1+\frac{1}{p},\ (p,r)\in[1,+\infty)\times\{1\}$. In our recent papers [33, 34], we established the ill-posedness for (1.1) in $B^{s}_{p,\infty}(\mathbb{R})$ with $s>\max\big{\{}1+1/p,3/2\big{\}}$ with $p\in[1,\infty]$ by proving the solution map to the CH equation starting from $u_{0}$ is discontinuous at $t=0$ in the metric of $B^{s}_{p,\infty}(\mathbb{R})$.
Assume that $u_{0}\in B^{s}_{p,r}$ with $s>1+\frac{1}{p},(p,r)\in[1,+\infty]\times[1,+\infty)$ or $s=1+\frac{1}{p},\ (p,r)\in[1,+\infty)\times\{1\}$, it is known that there exists a solution $u\in\mathcal{C}([0,T];B^{s}_{p,r})$ for the Camassa-Holm solution. Naturally, we may wonder whether or not the solution $u$ can belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with some $\alpha\in(0,1)$. Therefore, we are interested in the following question:
@@
$u_{0}\in B^{s}_{p,r}\;\Rightarrow\;\exists 1\;u\in\mathcal{C}([0,T];B^{s}_{p,r})\;\overset{?}{\Rightarrow}\;\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})\quad\text{with}\;\alpha\in(0,1).$
@@
If the initial data $u_{0}$ have more regularity such that $u_{0}\in B^{s^{\prime}}_{p,r}$ for some $s^{\prime}>s$, by the interpolation argument, we can deduce that $u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with $\alpha=s^{\prime}-s$. In this paper, we will show that there exits initial data $u_{0}\in B^{s}_{p,r}$ such that the corresponding solution of the Camassa-Holm solution can not belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with any $\alpha\in(0,1)$. Now let us state our main result of this paper.
@@
[…]
For any $\alpha\in(0,1)$, there exits $u_{0}\in B^{s}_{p,r}(\mathbb{R})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})\in\mathcal{C}([0,T];B^{s}_{p,r})$ of the Cauchy problem (1.2) satisfies
@@
Theorem 1.1 also holds for the following b-family equation (see [25, 26, 15, 27, 28] etc.):
$\begin{cases}\partial_{t}u+uu_{x}=-\partial_{x}(1-\partial^{2}_{x})^{-1}\big{(}\frac{b}{2}u^{2}+\frac{3-b}{2}u^{2}_{x}\big{)},&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R},\\ u(0,x)=u_{0}(x),&\quad x\in\mathbb{R}.\end{cases}$ (1.5)
It should be mentioned that the Camassa-Holm equation corresponds to $b=2$ and Degasperis-Procesi equation corresponds to $b=3$ [35].
@@
We summary the local well-posedness/ill-posdeness results of the Camassa-Holm equation in the Besov spaces. This can be seen clearly from the Table below.
@@
[13, 14, 32] $s>\max\big{\{}1+\frac{1}{p},\frac{3}{2}\big{\}},(p,r)\in[1,\infty]\times[1,\infty)$ and $s=\frac{3}{2},p=2,r=1$ LWP
[17] $s=1+\frac{1}{p},(p,r)\in[1,\infty]\times(1,\infty]$ Norm inflation
@@
[33, 34] $s>\max\big{\{}1+\frac{1}{p},\frac{3}{2}\big{\}},p\in[1,\infty],r=\infty$ Discontinuos
[40] $s>1+\frac{1}{p},p\in[1,+\infty],r\in[1,+\infty)$ and $s=1+\frac{1}{p},p\in[1,+\infty),r=1$ LWP
$s>1+\frac{1}{p},p\in[1,+\infty],r\in[1,+\infty)$ and $s=1+\frac{1}{p},p\in[1,+\infty),r=1$ Theorem 1.1
@@
The Cauchy problem for the Novikov equation reads as (see [19, 20, 29] etc.)
$\begin{cases}v_{t}+v^{2}v_{x}=-(1-\partial^{2}_{x})^{-1}\Big{(}\frac{1}{2}v_{x}^{3}+\partial_{x}\big{(}\frac{3}{2}vv^{2}_{x}+v^{3}\big{)}\Big{)}=:\mathbf{Q}(v),\\ v(0,x)=v_{0}.\end{cases}$ (1.6)
Then, we have the following
Theorem 1.2.
Assume that $(s,p,r)$ satisfies that (1.4). For any $\alpha\in(0,1)$, there exits $v_{0}\in B^{s}_{p,r}(\mathbb{R})$ such that the data-to-solution map $v_{0}\mapsto\mathbf{S}_{t}(v_{0})\in\mathcal{C}([0,T];B^{s}_{p,r})$ of the Cauchy problem (1.6) satisfies
$\limsup_{t\to 0^{+}}\frac{\|\mathbf{S}_{t}(v_{0})-v_{0}\|_{B^{s}_{p,r}}}{t^{\alpha}}=+\infty.$
@@
Following out method, Theorem 1.1 and Theorem 1.2 also hold for the Torus case. We omit the details and leave to the interesting readers.
@@ 2 Preliminaries
Notation  The notation $A\lesssim B$ (resp., $A\gtrsim B$) means that there exists a harmless positive constant $c$ such that $A\leq cB$ (resp., $A\geq cB$). Given a Banach space $X$, we denote its norm by $\|\cdot\|_{X}$. For $I\subset\mathbb{R}$, we denote by $\mathcal{C}(I;X)$ the set of continuous functions on $I$ with values in $X$. Sometimes we will denote $L^{p}(0,T;X)$ by $L_{T}^{p}X$.
Let us recall that for all $f\in\mathcal{S}^{\prime}$, the Fourier transform $\widehat{f}$, is defined by
$(\mathcal{F}f)(\xi)=\widehat{f}(\xi)=\int_{\mathbb{R}}e^{-\mathrm{i}x\xi}f(x)\mathrm{d}x\quad\text{for any}\;\xi\in\mathbb{R}.$
Next, we will recall some facts about the Littlewood-Paley decomposition and the nonhomogeneous Besov spaces (see [1] for more details). Choose a radial, non-negative, smooth function $\vartheta:\mathbb{R}\mapsto[0,1]$ such that ${\rm{supp}}\,\vartheta\subset B(0,4/3)$ and $\vartheta(\xi)\equiv 1$ for $|\xi|\leq 3/4$. Setting $\varphi(\xi):=\vartheta(\xi/2)-\vartheta(\xi)$, then we deduce that $\varphi$ has the following properties
  • ${\rm{supp}}\;\varphi\subset\left\{\xi\in\mathbb{R}:3/4\leq|\xi|\leq 8/3\right\}$;
  • $\varphi(\xi)\equiv 1$ for $24/3\leq|\xi|\leq 3/$ 2;
  • $\vartheta(\xi)+\sum_{j\geq 0}\varphi(2^{-j}\xi)=1$ for any $\xi\in\mathbb{R}$.
@@
$\Delta_{j}u=$ $0,$ if $j\leq-2$;
$\Delta_{j}u=$ $\chi(D)u=\mathcal{F}^{-1}(\chi\mathcal{F}u),$ if $j=-1$;
$\Delta_{j}u=$ $\varphi(2^{-j}D)u=\mathcal{F}^{-1}\big{(}\varphi(2^{-j}\cdot)\mathcal{F}u\big{)},$ if $j\geq 0$.
The inhomogeneous low-frequency cut-off operator $S_{j}$ is defined by
$S_{j}u=\sum_{q=-1}^{j-1}{\Delta}_{q}u.$
@@
where
$\|f\|_{B^{s}_{p,r}(\mathbb{R})}=$ $\left(\sum_{j\geq-1}2^{sjr}\|\Delta_{j}f\|^{r}_{L^{p}(\mathbb{R})}\right)^{\frac{1}{r}},$ if $1\leq r<\infty$,
$\|f\|_{B^{s}_{p,r}(\mathbb{R})}=$ $\sup_{j\geq-1}2^{sj}\|\Delta_{j}f\|_{L^{p}(\mathbb{R})},$ if $r=\infty$.
@@
$uv={T}_{u}v+{T}_{v}u+{R}(u,v)\quad\text{with}$
${T}_{u}v=\sum_{j\geq-1}{S}_{j-1}u{\Delta}_{j}v\quad\mbox{and}\quad{R}(u,v)=\sum_{|j-k|\leq 1}{\Delta}_{j}u{\Delta}_{k}v.$
@@
$\left\|T_{u}v\right\|_{B_{p,r}^{s}\left(\mathbb{R}^{d}\right)}\leq C^{|s|+1}\|u\|_{L^{\infty}\left(\mathbb{R}^{d}\right)}\|v\|_{B_{p,r}^{s}\left(\mathbb{R}^{d}\right)},$
$\left\|T_{u}v\right\|_{B_{p,r}^{s+t}\left(\mathbb{R}^{d}\right)}\leq\frac{C^{|s+t|+1}}{-t}\|u\|_{B_{\infty,r_{1}}^{t}\left(\mathbb{R}^{d}\right)}\|v\|_{B_{p,r_{2}}^{s}\left(\mathbb{R}^{d}\right)},$
$\|R(u,v)\|_{B_{p,r}^{s_{1}+s_{2}}\left(\mathbb{R}^{d}\right)}\leq\frac{C^{\left|s_{1}+s_{2}\right|+1}}{s_{1}+s_{2}}\|u\|_{B_{p_{1},r_{1}}^{s_{1}}\left(\mathbb{R}^{d}\right)}\|v\|_{B_{p_{2},r_{2}}^{s_{2}}\left(\mathbb{R}^{d}\right)}.$
@@
$\|\mathbf{P}(u)-\mathbf{P}(v)\|_{B_{p,r}^{s-1}}$ $\lesssim\|u-v\|_{B_{p,r}^{s-1}}\big{(}\|u\|_{B_{p,r}^{s}}+\|v\|_{B_{p,r}^{s}}\big{)}.$ (2.7)
@@
For $(p,r)\in[1,\infty]^{2}$, $s_{1}\neq s_{2}$ and $\theta\in(0,1)$, the following inequality holds
$\|u\|_{{B}_{p,r}^{\theta s_{1}+(1-\theta)s_{2}}}\leq\|u\|^{\theta}_{{B}_{p,r}^{s_{1}}}\|u\|^{1-\theta}_{{B}_{p,r}^{s_{2}}}.$
Lemma 2.5 (see [1]).
For $1\leq p\leq\infty$ and $s>0$. There exists a constant $C$, depending continuously on $p$ and $s$, we have
$\left\|2^{js}\left\|[\Delta_{j},v]\partial_{x}f\right\|_{L^{p}}\right\|_{\ell^{\infty}}\leq C\big{(}\|\partial_{x}v\|_{L^{\infty}}\|f\|_{B_{p,\infty}^{s}}+\|\partial_{x}f\|_{L^{\infty}}\|\partial_{x}v\|_{B_{p,\infty}^{s-1}}\big{)},$
where we denote the standard commutator $[\Delta_{j},v]\partial_{x}f=\Delta_{j}(v\partial_{x}f)-v\Delta_{j}\partial_{x}f$.
3 Proof of Main Theorems

### 14 Mar 2026 v2

# Nonexistence of Hölder continuous solution for the Camassa-Holm equation in Besov spaces

@@
Abstract: In the paper, we show that the continuity of the solution can not be improved to the Hölder continuity. Precisely speaking, the solution of the Camassa-Holm equation belongs to $\mathcal{C}([0,T];B^{s}_{p,r})$ but not to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with any $\alpha\in(0,1)$. To the best of our knowledge, our work is the first one addressing the issue on the failure of Hölder continuous in time of solution to the classical Camassa-Holm equation. As a by-product, we establish the ill-posedness for the Camassa-Holm equation in $B^{s}_{p,\infty}(\mathbb{R})$ with $s>\max\big\{1+1/p,3/2\big\}$ with $p\in[1,\infty]$ by proving the solution map to the Camassa-Holm equation starting from $u_{0}$ is discontinuous at $t=0$ in $B^{s}_{p,\infty}(\mathbb{R})$.
Keywords: Camassa-Holm equation, Hölder regularity, Besov spaces.
MSC (2010): 35Q35; 35A01.
@@ 1 Introduction
In this paper, we are concerned with the Cauchy problem for the classical Camassa-Holm (CH) equation
@@
Here the scalar function $u=u(t,x)$ stands for the fluid velocity at time $t\geq 0$ in the $x$ direction. Setting $\Lambda^{-2}=(1-\partial^{2}_{xx})^{-1}$, then $\Lambda^{-2}f=G*f$ where $G(x)=\frac{1}{2}e^{-|x|}$ is the kernel of the operator $\Lambda^{-2}$. Thus, we can transform the CH equation (1.1) equivalently into the following transport type equation
@@
$\begin{cases}\partial_{t}u+u\partial_{x}u=\mathbf{P}(u):=-\partial_{x}\Lambda^{-2}\left(u^{2}+\frac{1}{2}(\partial_{x}u)^{2}\right),\;&(x,t)\in\mathbb{R}\times\mathbb{R}^{+},\\ u(x,t=0)=u_{0}(x),\;&x\in\mathbb{R}.\end{cases}$ (1.2)
@@
The CH equation (1.1) was first proposed in the context of hereditary symmetries studied in [9] and then was derived explicitly as a water wave equation by Camassa-Holm [2]. Many aspects of the mathematical beauty of the CH equation have been exposed over the last two decades. Particularly, (1.1) is completely integrable [2, 5] with a bi-Hamiltonian structure [4, 9] and infinitely many conservation laws [2, 9]. Also, it admits exact peaked soliton solutions (peakons) of the form $u(x,t)=ce^{-|x-ct|}\;(c>0)$, which are orbitally stable [6]. Another remarkable feature of the CH equation is the wave breaking phenomena: the solution remains bounded while its slope becomes unbounded in finite time [3]. After the CH equation was derived physically in the context of water waves, there is a large amount of literature devoted to studying the well-posedness of the Cauchy problem (1.1) in Besov spaces. Danchin [7, 8] proved the local existence and uniqueness of strong solutions to (1.1) with initial data in $B^{s}_{p,r}$ if $(p,r)\in[1,\infty]\times[1,\infty),s>\max\big\{1+1/p,3/2\big\}$ and $B^{3/2}_{2,1}$. Meanwhile, he [7] only obtained the continuity of the solution map of (1.1) with respect to initial data in the space $\mathcal{C}([0,T];B^{s^{\prime}}_{p,r})$ with any $s^{\prime}<s$. Li-Yin [13] proved the continuity of the solution map of (1.1) with respect to initial data in the space $\mathcal{C}([0,T];B^{s}_{p,r})$ with $r<\infty$. In particular, they [13] proved that the solution map of (1.1) is weak continuous with respect to initial data $u_{0}\in B^{s}_{p,\infty}$. For the endpoints, Danchin [8] obtained that the data-to-solution map is not continuous by using peakon solution, which implies the ill-posedness of (1.1) in $B^{3/2}_{2,\infty}$. Guo-Liu-Molinet-Yin [10] showed the ill-posedness of (1.1) in $B_{p,r}^{1+1/p}(\mathbb{R}\;\text{or}\;\mathbb{T})$ with $(p,r)\in[1,\infty]\times(1,\infty]$ by proving the norm inflation. Guo-Ye-Yin [11] and Li-Yu-Zhu [14, 15] obtained the ill-posedness for the CH equation and the Novikov equation in $B^{1}_{\infty,1}$ by proving the norm inflation, respectively. Ye-Yin-Guo [16] obtained the local well-posedness for the Camassa-Holm type equation in $B^{s}_{p,r}$ with $s>1+\frac{1}{p},(p,r)\in[1,+\infty]\times[1,+\infty)$ or $s=1+\frac{1}{p},\ (p,r)\in[1,+\infty)\times\{1\}$. Assume that $u_{0}\in B^{s}_{p,r}$ with $s>1+\frac{1}{p},(p,r)\in[1,+\infty]\times[1,+\infty)$ or $s=1+\frac{1}{p},\ (p,r)\in[1,+\infty)\times\{1\}$, it is known that there exists a solution $u\in\mathcal{C}([0,T];B^{s}_{p,r})$ for the Camassa-Holm equation. Naturally, we may wonder whether or not the solution $u$ can belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with some $\alpha\in(0,1)$. Therefore, we are interested in the following Question:
@@
$u_{0}\in B^{s}_{p,r}\;\Rightarrow\;\exists 1\;u\in\mathcal{C}([0,T];B^{s}_{p,r})\;\overset{?}{\Rightarrow}\;u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})\quad\text{with}\;\alpha\in(0,1).$
@@
Assume that the initial data $u_{0}$ has more regularity such that $u_{0}\in B^{s^{\prime}}_{p,r}$ for some $s^{\prime}>s$, by the interpolation argument, we can deduce that $u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with $\alpha=s^{\prime}-s$. In this paper, we will show that that there exists initial data $u_{0}\in B^{s}_{p,r}$ such that the corresponding solution of the Camassa-Holm equation can not belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with any $\alpha\in(0,1)$. Now let us state our main result of this paper.
@@
$s>1+\frac{1}{p},\ (p,r)\in[1,\infty]\times[1,\infty)\quad\text{or}\quad s=1+\frac{1}{p},\ (p,r)\in[1,\infty)\times\{1\}.$ (1.3)
For any $\alpha\in(0,1)$, there exists $u_{0}\in B^{s}_{p,r}(\mathbb{R})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})\in\mathcal{C}([0,T];B^{s}_{p,r})$ of the Cauchy problem (1.2) satisfies
@@
In our recent papers [12], we established the ill-posedness for the Camassa-Holm equation in $B^{s}_{p,\infty}(\mathbb{R})$ with $s>\max\big\{1+1/p,3/2\big\}$ with $p\in[1,\infty]$ by proving the solution map to the CH equation starting from $u_{0}$ is discontinuous at $t=0$ in the metric of $B^{s}_{p,\infty}(\mathbb{R})$. As a by-product of Theorem 1.1, we have
Corollary 1.1.
Assume that $(s,p,r)$ satisfies that $s>1+\frac{1}{p}$ with $p\in[1,\infty]$. There exits $u_{0}\in B^{s}_{p,\infty}(\mathbb{R})$ and a positive constant $\varepsilon_{0}$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.2) satisfies
$\limsup_{t\to 0^{+}}\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,\infty}}\geq\varepsilon_{0}.$
@@
Compared with the previous result in [12], Corollary 1.1 is new since the regularity have been enlarged. In fact, based on some new observations and subtle estimations, we can deal with the case when $1+\frac{1}{p}<s\leq\frac{3}{2}$.
@@
We summary the local well-posedness/ill-posdeness results of the Camassa-Holm equation in Besov spaces. This can be seen clearly from the Table 1.
@@
[7, 8, 13] $s>\max\left\{1+1/p,3/2\right\},(p,r)\in[1,\infty]\times[1,\infty)$ $s=3/2,p=2,r=1$ LWP
[10] $s=1+1/p,(p,r)\in[1,\infty]\times(1,\infty]$ Norm inflation
@@
[16] $s>1+1/p,\ (p,r)\in[1,\infty]\times[1,\infty)$ $s=1+1/p,\ (p,r)\in[1,\infty)\times\{1\}$ LWP
Theorem 1.1 $s>1+1/p,\ (p,r)\in[1,\infty]\times[1,\infty)$ $s=1+1/p,\ (p,r)\in[1,\infty)\times\{1\}$ No Hölder continuos
Corollary 1.1 $s>1+1/p,\ p\in[1,\infty],r=\infty$ Discontinuos
@@
Theorem 1.1 is true for the Torus case and also holds for the b-family equation
$\begin{cases}\partial_{t}u+uu_{x}=-\partial_{x}(1-\partial^{2}_{x})^{-1}\left(\frac{b}{2}u^{2}+\frac{3-b}{2}u^{2}_{x}\right),&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R},\\ u(0,x)=u_{0}(x),&\quad x\in\mathbb{R},\end{cases}$
and the Novikov equation
$\begin{cases}\partial_{t}v+v^{2}v_{x}=-(1-\partial^{2}_{x})^{-1}\left(\frac{1}{2}v_{x}^{3}+\partial_{x}\left(\frac{3}{2}vv^{2}_{x}+v^{3}\right)\right),\\ v(0,x)=v_{0}(x).\end{cases}$
@@ 2 Preliminaries
Notation  The notation $A\lesssim B$ (resp., $A\gtrsim B$) means that there exists a harmless positive constant $c$ such that $A\leq cB$ (resp., $A\geq cB$). Given a Banach space $X$, we denote its norm by $\|\cdot\|_{X}$. For $I\subset\mathbb{R}$, we denote by $\mathcal{C}(I;X)$ the set of continuous functions on $I$ with values in $X$. Sometimes we will denote $L^{p}(0,T;X)$ by $L_{T}^{p}X$. Next, we will recall some facts about the Littlewood-Paley decomposition and the nonhomogeneous Besov spaces (see [1] for more details). Choose a radial, non-negative, smooth function $\vartheta:\mathbb{R}\mapsto[0,1]$ such that ${\rm{supp}}\,\vartheta\subset B(0,4/3)$ and $\vartheta(\xi)\equiv 1$ for $|\xi|\leq 3/4$. Setting $\varphi(\xi):=\vartheta(\xi/2)-\vartheta(\xi)$, then we deduce that $\varphi$ has the following properties ${\rm{supp}}\;\varphi\subset\left\{\xi\in\mathbb{R}:3/4\leq|\xi|\leq 8/3\right\}$ and $\varphi(\xi)\equiv 1$ for $4/3\leq|\xi|\leq 3/2$.
@@
$\Delta_{j}u=0,\;\text{if}\;j\leq-2;\quad\Delta_{-1}u=\vartheta(D)u;\quad\Delta_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\geq 0,$
The inhomogeneous low-frequency cut-off operator $S_{j}$ is defined by $S_{j}u=\sum_{q=-1}^{j-1}{\Delta}_{q}u.$
@@
where $\|f\|_{B^{s}_{p,\infty}(\mathbb{R})}=\sup_{j\geq-1}2^{sj}\|\Delta_{j}f\|_{L^{p}(\mathbb{R})}$ and for $1\leq r<\infty$,
$\|f\|_{B^{s}_{p,r}(\mathbb{R})}=\left(\sum_{j\geq-1}2^{sjr}\|\Delta_{j}f\|^{r}_{L^{p}(\mathbb{R})}\right)^{1/r}.$
@@
$uv={T}_{u}v+{T}_{v}u+{R}(u,v)\quad\text{with}\quad{T}_{u}v=\sum_{j\geq-1}{S}_{j-1}u{\Delta}_{j}v\quad\mbox{and}\quad{R}(u,v)=\sum_{|j-k|\leq 1}{\Delta}_{j}u{\Delta}_{k}v.$
@@
$\left\|T_{u}v\right\|_{B_{p,r}^{s}\left(\mathbb{R}\right)}\leq C^{|s|+1}\|u\|_{L^{\infty}\left(\mathbb{R}\right)}\|v\|_{B_{p,r}^{s}\left(\mathbb{R}\right)},$
$\left\|T_{u}v\right\|_{B_{p,r}^{s+t}\left(\mathbb{R}\right)}\leq\frac{C^{|s+t|+1}}{-t}\|u\|_{B_{\infty,r_{1}}^{t}\left(\mathbb{R}\right)}\|v\|_{B_{p,r_{2}}^{s}\left(\mathbb{R}\right)},$
$\|R(u,v)\|_{B_{p,r}^{s_{1}+s_{2}}\left(\mathbb{R}\right)}\leq\frac{C^{\left|s_{1}+s_{2}\right|+1}}{s_{1}+s_{2}}\|u\|_{B_{p_{1},r_{1}}^{s_{1}}\left(\mathbb{R}\right)}\|v\|_{B_{p_{2},r_{2}}^{s_{2}}\left(\mathbb{R}\right)}.$
@@
$\|\mathbf{P}(u)-\mathbf{P}(v)\|_{B_{p,r}^{s-1}}$ $\leq C\|u-v\|_{B_{p,r}^{s-1}}\left(\|u\|_{B_{p,r}^{s}}+\|v\|_{B_{p,r}^{s}}\right).$
@@
Let $p\in[1,\infty]$ and $s>0$. There exists a constant $C$, depending continuously on $p$ and $s$, such that
$\left\|2^{js}\|[\Delta_{j},v]\partial_{x}f\|_{L^{p}}\right\|_{\ell^{\infty}}\leq C\left(\|\partial_{x}v\|_{L^{\infty}}\|f\|_{B_{p,\infty}^{s}}+\|\partial_{x}f\|_{L^{\infty}}\|\partial_{x}v\|_{B_{p,\infty}^{s-1}}\right),$
where we denote the standard commutator $\Delta_{j},v]\partial_{x}f=\Delta_{j}(v\partial_{x}f)-v\Delta_{j}\partial_{x}f$.
3 Proof of Theorem 1.1